# A complete proof of the geometric Bombieri–Lang conjecture for ramified covers of abelian varieties

### v1

@@ 3.3 A positivity lemma
Consider the composition ${\mathcal{Z}}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$, which is surjective by our assumption. By Raynaud–Gruson’s flattening theorem (cf. [RG, Théomème 5.2.2]), there is a proper birational map $g\colon T^{\prime}\to T$ such that the strict transform $\mathcal{Z}$$^{\prime}$ of $\mathcal{Z}$} under $h$ is flat over $T^{\prime}$. Let $g_{{\mathcal{Z}}}\colon{\mathcal{Z}}^{\prime}\to{\mathcal{Z}}$ be the natural morphism. For each $n$, take a point $t^{\prime}_{n}\in T^{\prime}(\mathbb{C})$ such that $g(t^{\prime}_{n})=t_{n}$. Then we can lift the section $(s_{n},t_{n})\colon B\to{\mathcal{Z}}$ to a section $B\to{\mathcal{Z}}^{\prime}$. View $\mathcal{Z}$$^{\prime}$ as a closed subscheme of ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$, we can denote this section by $(s_{n},t^{\prime}_{n})$. As $T^{\prime}(\mathbb{C})$ is compact, after taking a subsequence, we can assume that $t^{\prime}_{n}$ tends to a point $t^{\prime}_{\infty}\in T^{\prime}(\mathbb{C})$. The flatness of ${\mathcal{Z}}^{\prime}\to T^{\prime}$ implies that all fibres ${\mathcal{Z}}^{\prime}_{t^{\prime}}$ has the same dimension. Since $f$ is not surjective, $Z$ is a proper closed subset of $A\times_{K}T_{K}$ and $\mathcal{Z}$} is a proper closed subset of ${\mathcal{A}}\times_{\mathbb{C}}T$. Therefore $\dim{\mathcal{Z}}^{\prime}=\dim{\mathcal{Z}}<\dim({\mathcal{A}}\times_{\mathbb{C}}T)=\dim({\mathcal{A}}\times_{\mathbb{C}}T^{\prime})$. This implies that ${\mathcal{Z}}^{\prime}_{t^{\prime}}$ is a proper closed subset of ${\mathcal{A}}\times\{t^{\prime}\}$ for all $t^{\prime}\in T^{\prime}$. In particular, ${\mathcal{Z}}^{\prime}_{t^{\prime}_{\infty}}$ is a proper closed subset of ${\mathcal{A}}\times\{t^{\prime}_{\infty}\}$. We then deduce that ${\mathcal{Z}}^{\prime}_{b,t^{\prime}_{\infty}}$ is a proper closed subset of ${\mathcal{A}}_{b}\times\{t^{\prime}_{\infty}\}$ for all but finitely many $b\in B(\mathbb{C})$. Here ${\mathcal{Z}}^{\prime}_{b,t^{\prime}_{\infty}}$ is the fibre over the point $(b,t^{\prime}_{\infty})$ under the natural map ${\mathcal{Z}}^{\prime}\to B\times_{\mathbb{C}}T^{\prime}$.
@@
The arguments of this non-surjective case essentially implies our main theorem when $X$ is a subvariety of abelian variety, which is known by [Fal3, Ray, Par, Bui, Hru].
@@
By [AK, Theorem 0.3], there is a generically finite projective surjective morphism $T^{\prime}\to T$ with $T^{\prime}$ smooth and a projective strict modification ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}$ such that $\mathcal{X}$$^{\prime}$ is normal and all fibres of the morphism ${\mathcal{X}}^{\prime}\to T^{\prime}$ are reduced. Moreover, ${\mathcal{X}}^{\prime}\to T^{\prime}$ is equidimensional. Here a strict modification ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}$ is a birational morphism from $\mathcal{X}$$^{\prime}$ to the Zariski closure of $T^{\prime}\times_{T}\eta$ in $T^{\prime}\times_{T}{\mathcal{X}}$, where $\eta$ is the generic point of $\mathcal{X}$}. Note that although $T^{\prime}\times_{T}{\mathcal{X}}$ may be reducible, $\mathcal{X}$$^{\prime}$ is irreducible.
@@
Let $T_{1}\to T^{\prime}$ be a finite morphism with $T_{1}$ smooth and (Brody) hyperbolic. In particular, $T_{1}$ does not contain rational curves. The existence of $T_{1}$ is essentially due to Hironaka; see [Gao, Lemma 2.1] for a proof (The author only proved the case of surfaces there, but the general situation is completely similar). Let ${\mathcal{X}}_{1}$ be the normalization of ${\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1}$. Then we have a generically finite morphism ${\mathcal{X}}_{1}\to{\mathcal{A}}\times_{\mathbb{C}}T_{1}$, which we denote by $f_{1}$. Let $X_{1}$ be the generic fibre of ${\mathcal{X}}_{1}\to B$ and let $R_{1}$ be the ramification divisor of the morphism $X_{1}\to A\times_{K}(T_{1})_{K}$. Denote by ${\mathcal{R}}_{1}$ the Zariski closure of $R_{1}$ in ${\mathcal{X}}_{1}$. For any $t_{1}\in T_{1}(\mathbb{C})$, let $t^{\prime}$ be the image of $t_{1}$ in $T^{\prime}(\mathbb{C})$. Using the same argument as in the previous paragraph, we can deduce that the étale locus of $({\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1})_{t_{1}}\to{\mathcal{A}}\times\{t_{1}\}$ is contained in the smooth locus of ${\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1}$. This means that the normalization ${\mathcal{X}}_{1}\to{\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1}$ is an isomorphism over this étale locus. In particular, we see that ${\mathcal{R}}_{1}$ is disjoint from the inverse image of this étale locus in ${\mathcal{X}}_{1}$. Moreover, as $({\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1})_{t_{1}}\cong{\mathcal{X}}^{\prime}_{t^{\prime}}$ is reduced and has the same dimension as ${\mathcal{A}}\times\{t_{1}\}$, the same argument as in the previous paragraph shows that the image of the non-étale locus of $({\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1})_{t_{1}}\to{\mathcal{A}}\times\{t_{1}\}$ in ${\mathcal{A}}\times\{t_{1}\}$ is a proper closed subset of ${\mathcal{A}}\times\{t_{1}\}$. Combining all these facts, we conclude that $f_{1}({\mathcal{R}}_{1})$ does not contain ${\mathcal{A}}\times\{t_{1}\}$ for any $t_{1}\in T_{1}$.
@@
Using Stein factorization for $f^{\prime}$, we get a birational morphism $\pi\colon{\mathcal{X}}^{\prime}\to{\mathcal{X}}^{\prime\prime}$ and a finite morphism $f^{\prime\prime}\colon{\mathcal{X}}^{\prime\prime}\to{\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$ such that $f^{\prime}=f^{\prime\prime}\circ\pi$. Moreover, $\mathcal{X}$$^{\prime\prime}$ is normal by [Laz, Example 2.1.15]. Let $x^{\prime\prime}_{n}$ be the image of $x^{\prime}_{n}$ in $\mathcal{X}$$^{\prime\prime}$. Let $X^{\prime\prime}$ be the generic fibre of ${\mathcal{X}}^{\prime\prime}\to B$. Let $R^{\prime\prime}$ be the ramification divisor of $X^{\prime\prime}\to A\times T^{\prime}_{K}$ and let $\mathcal{R}$$^{\prime\prime}$ be the Zariski closure of $R^{\prime\prime}$ in $\mathcal{X}$$^{\prime\prime}$.
@@
This means that we assume the original data satisfies the following properties:
  1. 1.
@@
Let $\pi\colon\widetilde{X}\to X$ be a resolution of singularity of $X$. Since $X$ is of general type, so is $\widetilde{X}$. By [BCHM, Theorem 1.1], the canonical ring $)\oplus_{m\geq 0}H^{0}(\widetilde{X},{\mathcal{O}}_{\widetilde{X}}(mK_{\widetilde{X}}$)) of $\widetilde{X}$ is finitely generated. Let $)X_{\rm can}={\rm Proj}\oplus_{m\geq 0}H^{0}(\widetilde{X},{\mathcal{O}}_{\widetilde{X}}(mK_{\widetilde{X}}$)) be the canonical model of $\widetilde{X}$ and let $\pi_{\rm can}\colon\widetilde{X}\dashrightarrow X_{\rm can}$ be the canonical rational map. Replace $\widetilde{X}$ by a resolution of singularity of $X_{\rm can}$ if necessary, we may assume that $\pi_{\rm can}$ is a morphism. By [Kol, Theorem 1.15], which is due to Reid, $X_{\rm can}$ is normal, projective, birational to $X$ and has canonical singularities. In particular, $X_{\rm can}$ has dlt singularities. Moreover, the canonical divisor $K_{X_{\rm can}}$ of $X_{\rm can}$ is $\mathbb{Q}$-Cartier and ample. Since $X$ does not contain rational curves, by a theorem of Hacon-Mckernan (cf. [HM, Corollary 1.7]), there is a birational morphism $\varpi\colon X_{\rm can}\to X$ extending the rational map $\pi\circ(\pi_{\rm can})^{-1}\colon X_{\rm can}\dashrightarrow X$. Consider the composition $X_{\rm can}\overset{\varpi}{\to}X\overset{f}{\to}A\times_{K}T_{K}$ which is generically finite. Let $R_{\rm can}$ be the ramification divisor of this morphism. By Hurwitz’s formula,
@@
For general point $t\in T_{K}$, $R_{\rm can}\cap(X_{\rm can})_{t}$ is of codimension 1 in $(X_{\rm can})_{t}$, which is an ample divisor on $(X_{\rm can})_{t}$ by the relative ampleness of $R_{\rm can}$ over $T$. Furthermore, the restriction map $(X_{\rm can})_{t}\to A\times\{t\}$ is generically finite for general $t$. Thus by Lemma 3.10, we see that $(f\circ\varpi)_{*}(R_{\rm can}\cap(X_{\rm can})_{t})$ is ample on $A\times\{t\}$ for general $t\in T_{K}$. Note that the support of $(f\circ\varpi)_{*}(R_{\rm can}\cap(X_{\rm can})_{t})$ is contained in $D\cap(A\times\{t\})$. Thus by remark 3.11, the restriction of ${\mathcal{O}}_{A\times T_{K}}(D)$ on $A\times\{t\}$ is ample for general $t$. Finally by [FC, Remark 1.10(b)], the restriction of ${\mathcal{O}}_{A\times T_{K}}(D)$ on $A\times\{t\}$ is ample for every $t\in T_{K}$. In other words, $D\cap(A\times\{t\})$ is ample on $A\times\{t\}$. ∎
@@
Since $K_{X_{\rm can}}$ is ample on $X_{\rm can}$, there exists a positive rational number $c_{1}>0$ such that $K_{X_{\rm can}}-c_{1}(f\circ\varpi)^{*}D_{\rm red}$ is also ample. By (3.4.1), we see that $R_{\rm can}-c_{1}(f\circ\varpi)^{*}D_{\rm red}$ is relatively ample over $T$. Fix an ample divisor $A_{T}$ on $T$ whose support does not contain $t_{\infty}$. Now by [Laz, Proposition 1.7.10], $R_{\rm can}-c_{1}(f\circ\varpi)^{*}D_{\rm red}+n(f\circ\varpi)^{*}A_{T}$ is ample on $X_{\rm can}$ for some positive integer $n$. Thus it is $\mathbb{Q}$-linearly equivalent to an effective $\mathbb{Q}$-Cartier divisor $E$ on $X_{\rm can}$ such that $(f\circ\varpi)({\rm Supp}(E))$ does not contain $A\times\{t_{\infty}\}$. Now we have
@@
$\widetilde{{\mathcal{R}}}-c_{1}(f\circ\pi)^{*}{\mathcal{D}}_{\rm red}\sim\widetilde{{\mathcal{E}}}-\widetilde{{\mathcal{V}}}_{T}-\widetilde{{\mathcal{V}}}_{B}.$ (3.4.2)
@@
Tangency
As we have mentioned in §1.2, a key idea in the proof of general case is a tangency observation. Now we give the rigorous statement.
Let $\widetilde{{\mathcal{R}}}_{0}\subseteq\widetilde{{\mathcal{R}}}$ be the union of all irreducible components of $\widetilde{{\mathcal{R}}}$ which dominates an irreducible component of $\mathcal{R}$}. Then we have the following proposition.
Proposition 3.16.
Proof.
Corollary 3.17.
Proof.
Final proof
Now we come to the final proof.

### v2

@@ 3.3 Step 1: algebraic-geometric reductions
Consider the composition ${\mathcal{Z}}\hookrightarrow{\mathcal{A}}\times_{\mathbb{C}}T\to T$, which is surjective by our assumption. By Raynaud–Gruson’s flattening theorem (cf. [RG71, Théomème 5.2.2]), there is a proper birational map $g\colon T^{\prime}\to T$ such that the strict transform $\mathcal{Z}$$^{\prime}$ of $\mathcal{Z}$} under $h$ is flat over $T^{\prime}$. Let $g_{{\mathcal{Z}}}\colon{\mathcal{Z}}^{\prime}\to{\mathcal{Z}}$ be the natural morphism. For each $n$, take a point $t^{\prime}_{n}\in T^{\prime}(\mathbb{C})$ such that $g(t^{\prime}_{n})=t_{n}$. Then we can lift the section $(s_{n},t_{n})\colon B\to{\mathcal{Z}}$ to a section $B\to{\mathcal{Z}}^{\prime}$. View $\mathcal{Z}$$^{\prime}$ as a closed subscheme of ${\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$, we can denote this section by $(s_{n},t^{\prime}_{n})$. As $T^{\prime}(\mathbb{C})$ is compact, after taking a subsequence, we can assume that $t^{\prime}_{n}$ tends to a point $t^{\prime}_{\infty}\in T^{\prime}(\mathbb{C})$. The flatness of ${\mathcal{Z}}^{\prime}\to T^{\prime}$ implies that all fibres ${\mathcal{Z}}^{\prime}_{t^{\prime}}$ has the same dimension. Since $f$ is not surjective, $Z$ is a proper closed subset of $A\times_{K}T_{K}$ and $\mathcal{Z}$} is a proper closed subset of ${\mathcal{A}}\times_{\mathbb{C}}T$. Therefore $\dim{\mathcal{Z}}^{\prime}=\dim{\mathcal{Z}}<\dim({\mathcal{A}}\times_{\mathbb{C}}T)=\dim({\mathcal{A}}\times_{\mathbb{C}}T^{\prime})$. This implies that ${\mathcal{Z}}^{\prime}_{t^{\prime}}$ is a proper closed subset of ${\mathcal{A}}\times\{t^{\prime}\}$ for all $t^{\prime}\in T^{\prime}$. In particular, ${\mathcal{Z}}^{\prime}_{t^{\prime}_{\infty}}$ is a proper closed subset of ${\mathcal{A}}\times\{t^{\prime}_{\infty}\}$. We then deduce that ${\mathcal{Z}}^{\prime}_{b,t^{\prime}_{\infty}}$ is a proper closed subset of ${\mathcal{A}}_{b}\times\{t^{\prime}_{\infty}\}$ for all but finitely many $b\in B(\mathbb{C})$. Here ${\mathcal{Z}}^{\prime}_{b,t^{\prime}_{\infty}}$ is the fibre over the point $(b,t^{\prime}_{\infty})$ under the natural map ${\mathcal{Z}}^{\prime}\to B\times_{\mathbb{C}}T^{\prime}$.
@@
The arguments of this non-surjective case essentially implies our main theorem when $X$ is a subvariety of abelian variety, which is known by [FAL94, RAY83, PAR90, BUI92, HRU96].
@@
By [AK00, Theorem 0.3], there is a generically finite projective surjective morphism $T^{\prime}\to T$ with $T^{\prime}$ smooth and a projective strict modification ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}$ such that $\mathcal{X}$$^{\prime}$ is normal and all fibres of the morphism ${\mathcal{X}}^{\prime}\to T^{\prime}$ are reduced. Moreover, ${\mathcal{X}}^{\prime}\to T^{\prime}$ is equidimensional. Here a strict modification ${\mathcal{X}}^{\prime}\to{\mathcal{X}}\times_{T}T^{\prime}$ is a birational morphism from $\mathcal{X}$$^{\prime}$ to the Zariski closure of $T^{\prime}\times_{T}\eta$ in $T^{\prime}\times_{T}{\mathcal{X}}$, where $\eta$ is the generic point of $\mathcal{X}$}. Note that although $T^{\prime}\times_{T}{\mathcal{X}}$ may be reducible, $\mathcal{X}$$^{\prime}$ is irreducible.
@@
Let $T_{1}\to T^{\prime}$ be a finite morphism with $T_{1}$ smooth and (Brody) hyperbolic. In particular, $T_{1}$ does not contain rational curves. The existence of $T_{1}$ is essentially due to Hironaka; see [14, Lemma 2.1] for a proof (The author only proved the case of surfaces there, but the general situation is completely similar). Let ${\mathcal{X}}_{1}$ be the normalization of ${\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1}$. Then we have a generically finite morphism ${\mathcal{X}}_{1}\to{\mathcal{A}}\times_{\mathbb{C}}T_{1}$, which we denote by $f_{1}$. Let $X_{1}$ be the generic fibre of ${\mathcal{X}}_{1}\to B$ and let $R_{1}$ be the ramification divisor of the morphism $X_{1}\to A\times_{K}(T_{1})_{K}$. Denote by ${\mathcal{R}}_{1}$ the Zariski closure of $R_{1}$ in ${\mathcal{X}}_{1}$. For any $t_{1}\in T_{1}(\mathbb{C})$, let $t^{\prime}$ be the image of $t_{1}$ in $T^{\prime}(\mathbb{C})$. Using the same argument as in the previous paragraph, we can deduce that the étale locus of $({\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1})_{t_{1}}\to{\mathcal{A}}\times\{t_{1}\}$ is contained in the smooth locus of ${\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1}$. This means that the normalization ${\mathcal{X}}_{1}\to{\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1}$ is an isomorphism over this étale locus. In particular, we see that ${\mathcal{R}}_{1}$ is disjoint from the inverse image of this étale locus in ${\mathcal{X}}_{1}$. Moreover, as $({\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1})_{t_{1}}\simeq{\mathcal{X}}^{\prime}_{t^{\prime}}$ is reduced and has the same dimension as ${\mathcal{A}}\times\{t_{1}\}$, the same argument as in the previous paragraph shows that the image of the non-étale locus of $({\mathcal{X}}^{\prime}\times_{T^{\prime}}T_{1})_{t_{1}}\to{\mathcal{A}}\times\{t_{1}\}$ in ${\mathcal{A}}\times\{t_{1}\}$ is a proper closed subset of ${\mathcal{A}}\times\{t_{1}\}$. Combining all these facts, we conclude that $f_{1}({\mathcal{R}}_{1})$ does not contain ${\mathcal{A}}\times\{t_{1}\}$ for any $t_{1}\in T_{1}$.
@@
Using Stein factorization for $f^{\prime}$, we get a birational morphism $\pi\colon{\mathcal{X}}^{\prime}\to{\mathcal{X}}^{\prime\prime}$ and a finite morphism $f^{\prime\prime}\colon{\mathcal{X}}^{\prime\prime}\to{\mathcal{A}}\times_{\mathbb{C}}T^{\prime}$ such that $f^{\prime}=f^{\prime\prime}\circ\pi$. Moreover, $\mathcal{X}$$^{\prime\prime}$ is normal by [LAZ04, Example 2.1.15]. Let $x^{\prime\prime}_{n}$ be the image of $x^{\prime}_{n}$ in $\mathcal{X}$$^{\prime\prime}$. Let $X^{\prime\prime}$ be the generic fibre of ${\mathcal{X}}^{\prime\prime}\to B$. Let $R^{\prime\prime}$ be the ramification divisor of $X^{\prime\prime}\to A\times T^{\prime}_{K}$ and let $\mathcal{R}$$^{\prime\prime}$ be the Zariski closure of $R^{\prime\prime}$ in $\mathcal{X}$$^{\prime\prime}$.
@@
This means that we assume the original data satisfy the following properties:
  1. (1)
@@
Let $\pi\colon\widetilde{X}\to X$ be a resolution of singularity of $X$. Since $X$ is of general type, so is $\widetilde{X}$. By [BCH+10, Corollary 1.1.1(3)], the canonical ring $)\oplus_{m\geq 0}H^{0}(\widetilde{X},{\mathcal{O}}_{\widetilde{X}}(mK_{\widetilde{X}}$)) of $\widetilde{X}$ is finitely generated. Let $)X_{\rm can}={\rm Proj}\oplus_{m\geq 0}H^{0}(\widetilde{X},{\mathcal{O}}_{\widetilde{X}}(mK_{\widetilde{X}}$)) be the canonical model of $\widetilde{X}$ and let $\pi_{\rm can}\colon\widetilde{X}\dashrightarrow X_{\rm can}$ be the canonical rational map. Replace $\widetilde{X}$ by a resolution of singularity of $X_{\rm can}$ if necessary, we may assume that $\pi_{\rm can}$ is a morphism. By [KOL13, Theorem 1.15], which is due to Reid, $X_{\rm can}$ is normal, projective, birational to $X$ and has canonical singularities. In particular, $X_{\rm can}$ has dlt singularities. Moreover, the canonical divisor $K_{X_{\rm can}}$ of $X_{\rm can}$ is $\mathbb{Q}$-Cartier and ample. Since $X$ does not contain rational curves, by a theorem of Hacon-Mckernan (cf. [HM07, Corollary 1.7]), there is a birational morphism $\varpi\colon X_{\rm can}\to X$ extending the rational map $\pi\circ(\pi_{\rm can})^{-1}\colon X_{\rm can}\dashrightarrow X$. Consider the composition $X_{\rm can}\overset{\varpi}{\to}X\overset{f}{\to}A\times_{K}T_{K}$ which is generically finite. Let $R_{\rm can}$ be the ramification divisor of this morphism. By Hurwitz’s formula,
@@
For general point $t\in T_{K}$, $R_{\rm can}\cap(X_{\rm can})_{t}$ is of codimension 1 in $(X_{\rm can})_{t}$, which is an ample divisor on $(X_{\rm can})_{t}$ by the relative ampleness of $R_{\rm can}$ over $T$. Furthermore, the restriction map $(X_{\rm can})_{t}\to A\times\{t\}$ is generically finite for general $t$. Thus by Lemma 3.10, we see that $(f\circ\varpi)_{*}(R_{\rm can}\cap(X_{\rm can})_{t})$ is ample on $A\times\{t\}$ for general $t\in T_{K}$. Note that the support of $(f\circ\varpi)_{*}(R_{\rm can}\cap(X_{\rm can})_{t})$ is contained in $D\cap(A\times\{t\})$. Thus by remark 3.11, the restriction of ${\mathcal{O}}_{A\times T_{K}}(D)$ on $A\times\{t\}$ is ample for general $t$. Finally by [FC90, Remark 1.10(b)], the restriction of ${\mathcal{O}}_{A\times T_{K}}(D)$ on $A\times\{t\}$ is ample for every $t\in T_{K}$. In other words, $D\cap(A\times\{t\})$ is ample on $A\times\{t\}$. ∎
@@
Since $K_{X_{\rm can}}$ is ample on $X_{\rm can}$, there exists a positive rational number $c_{1}>0$ such that $K_{X_{\rm can}}-c_{1}(f\circ\varpi)^{*}D_{\rm red}$ is also ample. By (3.3.1), we see that $R_{\rm can}-c_{1}(f\circ\varpi)^{*}D_{\rm red}$ is relatively ample over $T$. Fix an ample divisor $A_{T}$ on $T$ whose support does not contain $t_{\infty}$. Now by [LAZ04, Proposition 1.7.10], $R_{\rm can}-c_{1}(f\circ\varpi)^{*}D_{\rm red}+n(f\circ\varpi)^{*}A_{T}$ is ample on $X_{\rm can}$ for some positive integer $n$. Thus it is $\mathbb{Q}$-linearly equivalent to an effective $\mathbb{Q}$-Cartier divisor $E$ on $X_{\rm can}$ such that $(f\circ\varpi)({\rm Supp}(E))$ does not contain $A\times\{t_{\infty}\}$. Now we have
@@
$\widetilde{{\mathcal{R}}}-c_{1}(f\circ\pi)^{*}{\mathcal{D}}_{\rm red}\sim\widetilde{{\mathcal{E}}}-\widetilde{{\mathcal{V}}}_{T}-\widetilde{{\mathcal{V}}}_{B},$ (3.3.2)
@@
As $\pi\colon\widetilde{X}\to X$ is birational and $\{x_{n}\}_{n\geq 1}$ is generic in $X$, after taking a subsequence, we may assume that $x_{n}$ can be uniquely lifted to a point $\tilde{x}_{n}\in\widetilde{X}(K)$ for all $n\geq 1$. Similarly as before, we often view $\tilde{x}_{n}$ as a section $\tilde{x}_{n}\colon B\to\widetilde{{\mathcal{X}}}$.
3.4 Step 2: lower bound of transversal intersections
Now we come to the first estimate in the counting argument. We will give an upper bound (see Lemma 3.16) on the number of tangency points between the sections $\{(s_{n},t_{n})\}_{n\geq 1}$ and $\widetilde{{\mathcal{D}}}_{\rm red}$, or equivalently, a lower bound on their transversal intersection numbers. The main ingredients in the estimate are Theorem 3.6(1) and Lemma 3.7. All the estimates in this subsection are performed on ${\mathcal{A}}\times_{\mathbb{C}}T$ and do not involve $\widetilde{{\mathcal{X}}}$.
Before estimating, we need to fix some notations. In order to present the proof in a logical order, part of these notations will only be used in the next subsection. Below we fix all the data and indicate which items will be needed later.
Choice of all data